If the incident photon has the wavelength eual to 2 Angstrom and it's angle of deflection is 90 degree, the kinetic energy of the recoil electron in the compton scattering ould be?
1
Expert's answer
2020-06-05T11:33:51-0400

Given, Incident wavelength of photon "\lambda =2\times10^{-10}m" ,Angle of deflection is "\theta =90^{\circ}" .

Now. we know that

"\lambda'-\lambda =\frac{h}{m_ec}(1-\cos(\theta)"

Thus,

"\lambda'=\lambda + \frac{h}{m_ec}=202.43\times 10^{-12}m"

Let, "K_e" is the energy of electron after collision.

As, energy is conserved in the collision ,

"\frac{hc}{\lambda}=\frac{hc}{\lambda'}+K_e\\
\implies K_e=hc\bigg(\frac{1}{\lambda}-\frac{1}{\lambda'}\bigg)\\
\implies K_e=1.98644568×10^{-25}\times 6.0020\times 10^7=11.9227955\times 10^{-18}J\\
\implies K_e=74.42eV"